### v1

# Universal Incremental Learning: Mitigating Confusion from Inter- and Intra-task Distribution Randomness

Abstract
Incremental learning (IL) aims to overcome catastrophic forgetting of previous tasks while learning new ones. Existing IL methods make strong assumptions that the incoming task type will either only increases new classes or domains (i.e. Class IL, Domain IL), or increase by a static scale in a class- and domain-agnostic manner (i.e. Versatile IL (VIL)), which greatly limit their applicability in the unpredictable and dynamic wild. In this work, we investigate Universal Incremental Learning (UIL), where a model neither knows which new classes or domains will increase along sequential tasks, nor the scale of the increments within each task. This uncertainty prevents the model from confidently learning knowledge from all task distributions and symmetrically focusing on the diverse knowledge within each task distribution. Consequently, UIL presents a more general and realistic IL scenario, making the model face confusion arising from inter-task and intra-task distribution randomness. To Mitigate both Confusion, we propose a simple yet effective framework for UIL, named MiCo. At the inter-task distribution level, we employ a multi-objective learning scheme to enforce accurate and deterministic predictions, and its effectiveness is further enhanced by a direction recalibration module that reduces conflicting gradients. Moreover, at the intra-task distribution level, we introduce a magnitude recalibration module to alleviate asymmetrical optimization towards imbalanced class distribution. Extensive experiments on three benchmarks demonstrate the effectiveness of our method, outperforming existing state-of-the-art methods in both the UIL scenario and the VIL scenario. Our code will be available at here.
Figure 1: Comparison of existing IL scenarios and our UIL scenario on a 6-class dataset with 4 domains. The horizontal axis represents class variation, while the vertical axis represents domain variation. Each incremental task is framed by a black box. UIL is a more general IL scenario than existing IL scenarios by simultaneously considering randomness in both the type and scale of increments, as the model neither knows which new classes or domains will increase across all task distributions, nor the number of new classes or domains within each task distribution.
1 Introduction
Incremental learning (IL) [29, 15, 38, 37, 33, 8, 40, 3] aims to continuously learn new knowledge over time. This paradigm is particularly crucial in the dynamic real-world scenario, as it allows a single model to incrementally learn all sequentially arrived tasks, without the need to respectively train task-specific models from scratch.
Based on the type of increments, mainstream IL scenarios are categorized into class-incremental learning (CIL)  [39, 38, 33] and domain-incremental learning (DIL) [37, 6], as illustrated in Fig. 1: 1) CIL assumes that each incremental task only introduces a disjoint class set with the same domains. 2) DIL only considers domain variation, where each incremental task introduces a different domain while sharing the same class set. However, both types simplify the IL scenario by concentrating on a single type of increment (i.e., either increasing new classes or domains), limiting their applicability. Actually, various types of increments across task distributions often arrive randomly in the unpredictable wild. For instance, the discovery of unknown classes frequently coincides with new domains due to varying weather conditions in realistic road scenes [22, 24]. More recently, Versatile Incremental Learning (VIL) [26] has been proposed to simulate the class- and domain-agnostic IL scenario without prior knowledge of which new classes or domains would increase, as illustrated in Fig. 1. However, VIL assumes that the scale of these increments (i.e., the number of new classes or domains) is static, which still significantly differs from dynamic reality where the scale of a new task is constantly changing. To relax these assumptions, it is necessary to explore a more general incremental learning scenario that simultaneously considers the randomness of incremental task types across all task distributions and the scale of increments within each task distribution. Such a scenario better reflects real-world conditions.
In this paper, we investigate a new IL scenario called Universal Incremental Learning (UIL), as illustrated in Fig. 1. In UIL, each incoming task has a random type of increments, which may involve new classes with previous domains, new domains with previous classes, or entirely new classes with new domains. Moreover, the scale of these increments is also random. The word “universal”indicates that UIL has no prior knowledge of both the type and scale of increments for each incoming task. In UIL, it poses two significant challenges for existing IL methods: 1) Since incremental types vary randomly across all task distributions, designing specific learnable modules tailored to a single type, as in existing IL methods [39, 38, 37], becomes ineffective. Moreover, when the model encounters a new incremental type, it will learn conflicting knowledge compared to what has already acquired in previous tasks, hindering accurate and deterministic predictions and thus causing catastrophic forgetting. We refer to this challenge as inter-task distribution confusion arising from random incremental types. 2) Given the random scale of increments within each task distribution, the class distribution evolves inherently imbalanced due to various emerging frequencies in the dynamic wild. Standard optimization algorithms [16, 14] calculate the average loss on all samples and update the parameters accordingly. However, a major challenge in training on imbalanced class distribution is asymmetrical optimization: gradient updates favor head classes with more samples, making the average gradient updates detrimental to tail classes with fewer samples. Consequently, tail classes easily suffer from under-fitting while head classes over-fit. We refer to this challenge as intra-task distribution confusion arising from random incremental scale.
To address the challenges of our UIL, we propose a simple yet effective method named MiCo, which individually Mitigating Confusion arising from inter- and intra-task distribution randomness. This method employs a multi-objective learning scheme and introduces direction- and magnitude-decoupled recalibration modules. Specifically, cross-entropy loss $\mathcal{L}_{ce}$ guides the model to distinguish correct labels by narrowing the distribution gap between prediction and ground truth, while entropy minimization loss $\mathcal{L}_{em}$ [35, 25] encourages the model to make deterministic predictions via explicitly reducing the prediction distribution entropy (PDE) of the samples, contributing to accurate and deterministic predictions regardless of how the incremental type changes. With the favor of reducing conflicting gradients (i.e. gradient cosine similarity (GCS) $\cos\theta<0$) in the direction recalibration module, the effectiveness of multi-objective learning scheme is further enhanced. In addition, asymmetrical optimization towards imbalanced class distribution is alleviated via the magnitude recalibration module, based on a high positive relationship between class distribution and gradient magnitude for each class. Overall, our main contributions are highlighted as follows:
  1. 1.
    We investigate a more general and realistic IL scenario than existing ones called Universal Incremental Learning (UIL), by simultaneously considering randomness in the type and scale of increments.
  2. 2.
    A simple yet effective framework named MiCo is proposed to mitigate confusion arising from inter- and intra-task distribution randomness in UIL, respectively.
  3. 3.
    Extensive experiments on three benchmarks demonstrate that our MiCo outperforms existing state-of-the-art methods in both the proposed UIL scenario and the existing VIL scenario.
2 Related Work
2.1 Incremental Learning
Incremental Learning (IL) faces the core challenge of overcoming catastrophic forgetting. This phenomenon refers to the model’s tendency to forget previously learned knowledge while acquiring new knowledge. To address this issue, numerous methods have been proposed. For example, rehearsal-based methods [21, 8, 12, 13] alleviate forgetting by replaying real or synthetic samples from previous tasks. These methods include experience replay [21, 8], which utilizes stored experiences from a memory buffer, and generative replay [12, 13], where a generative model creates synthetic data to simulate previous knowledge. Regularization-based methods maintain a relative balance between stability and plasticity during the learning of new tasks. Representative methods [26, 42, 15] involve designing the regularization loss that protects important parameters, ensuring that the model retains crucial information from earlier tasks while adapting to new ones. With the rise of tuning pre-trained models [17, 11, 43], prompt-based methods [39, 33, 38] have emerged as a promising alternative for IL. By injecting a small set of learnable parameters, these methods prompt the model to acquire additional instructions, enabling it to learn both task-specific and task-agnostic knowledge throughout a sequence of tasks.
2.2 Entropy Minimization
Entropy minimization, serving as a crucial regularizer, has been widely studied to improve generalization against domain shift in domain adaptation [31, 30] and test-time adaptation [35, 41]. Moreover, in the context of semi-supervised learning [1], entropy minimization is employed to enhance the robustness of models against random appearance variation. These methods demonstrate that reducing entropy can achieve excellent performance on out-of-distribution and unknown data. Our UIL confuses the model with random incremental types (i.e. ever-varying classes and domains) across all task distributions, so entropy minimization is an explicit solution to mitigate this confusion.
2.3 Optimization for Incremental Learning
Optimization-based methods for IL [20, 2, 32, 5, 8, 36] manipulate gradient direction to overcome catastrophic forgetting. For example, existing methods [20, 2] constrain the gradient direction of new tasks to align with the rehearsals’ direction, ensuring consistency in gradient spaces. GPM [32] and OGD [5] share similarities in preserving the gradient direction of the current task relative to previous tasks through explicit gradient projection and rectification, respectively. DGR [8] focuses on rectifying the gradient magnitude for each class while constraining parameter updates by distillation rehearsal’s distribution to preserve the learned knowledge. However, these optimization-based methods struggle in random task distributions. Constraining gradient direction to previous tasks causes significant forgetting due to inherent direction variation in a dynamically varying IL scenario [26].
3 Universal Incremental Learning
3.1 Problem Formulation and Definition
Problem Formulation. We formally introduce the definition of Universal Incremental Learning (UIL). Let $\mathcal{D}=\{D_{1},D_{2},\cdots,D_{T}\}$ denote a sequence of $T$ disjoint tasks. Each task $D_{t}=\{\mathcal{X}_{t},\mathcal{Y}_{t}\}$ contains $n_{t}$ images $\mathcal{X}_{t}=\{x_{1},x_{2},\cdots,x_{n_{t}}\}$ and the corresponding labels $\mathcal{Y}_{t}=\{y_{1},y_{2},\cdots,y_{n_{t}}\}$. For $i\in[1,2,...,n_{t}]$, each label $y_{i}=(c_{i},m_{i})$ is annotated with the class label $c_{i}\in\mathcal{C}_{t}$ and the domain label $m_{i}\in\mathcal{M}_{t}$. Formally, for $t,t^{\prime}\in[1,2,\cdots,T]$, disjoint tasks without overlapping samples can be expressed as $\mathcal{X}_{t}\cap\mathcal{X}_{t^{\prime}}=\emptyset$ and $\mathcal{Y}_{t}\cap\mathcal{Y}_{t^{\prime}}=\emptyset$ for $t\neq t^{\prime}$. UIL aims to simultaneously consider randomness from the type and scale of increments. Therefore, the size of the class set $\mathcal{C}_{t}$ and the domain set $\mathcal{M}_{t}$, represented by $||\mathcal{C}_{t}||$ and $||\mathcal{M}_{t}||$, are random. The model $f_{T}^{\theta}$ parameterized by $\theta_{T}$ at task $T$ is only trained on training set $D_{train}=D_{T}$ but is evaluated on test set $D_{test}=\{D_{1},D_{2},\cdots,D_{T}\}$ to achieve overall performance across $T$ tasks. To achieve this, the model is trained using the following objective, as shown in Eq. 1:
$\mathcal{L}_{\theta}=\sum\limits_{t=1}^{T}\mathbb{E}_{x,y\sim\mathcal{X}_{t},\mathcal{Y}_{t}}L(f^{\theta}_{t}(x),y),$ (1)
where $L$ is a common cross-entropy loss for classification.
In terms of $||\mathcal{C}_{t}||$ and $||\mathcal{M}_{t}||$ at task $t$, the universality of our proposed UIL is evident in the fact that the VIL scenario is actually a special case of the UIL scenario when $||\mathcal{C}_{t}||$ = $C$ and $||\mathcal{M}_{t}||=1$, where $C$ is a static constant.
Definition 3.1 (Prediction Distribution Entropy, PDE). At task $t$, we denote $\hat{y}_{i,c}$ as the logits of the $i$-th sample $x_{i}$ at $c$-th class index, predicted by the model $f^{\theta}_{t}$. $p(\hat{y}_{i,c})$ is corresponding probability after softmax operation. The PDE of $x_{i}$ is defined in Eq. 2:
$H(y_{i})=-\sum\limits_{c=0}^{||\mathcal{C}_{t}||}p(\hat{y}_{i,c})\log p(\hat{y}_{i,c}).$ (2)
As shown in Eq. 3, the output probability $p(\hat{y}_{i})$ follows a uniform distribution, indicating complete confusion in its ability to discriminate among $\mathcal{C}_{t}$, with highest value $\log||\mathcal{C}_{t}||$. Conversely, when the output probability $p(\hat{y}_{i})$ follows a one-hot distribution, reflecting a fully deterministic prediction, with the lowest value 0.
$0\leq H(y_{i})\leq\log||\mathcal{C}_{t}||.$ (3)
Definition 3.2 (Gradient Cosine Similarity, GCS). Denote $\theta_{u,v}$ as the angle between two gradients $g_{u}$ and $g_{v}$, GCS is denoted as $\cos\theta_{u,v}$ between $g_{u}$ and $g_{v}$, as shown in Eq. 4:
$\cos\theta_{u,v}=\frac{g_{u}}{||g_{u}||}\cdot\frac{g_{v}}{||g_{v}||},$ (4)
where the conflicting gradients occur when $\cos\theta_{u,v}<0$.
(a) Distribution comparison of prediction distribution entropy (PDE) between UIL and VIL scenarios. Our UIL causes more confusion to the model than VIL with larger entropy mean and variance (i.e. 2.19 vs 1.54 and 2.25 vs 1.78), due to the smaller proportion of samples in the low-entropy intervals and the larger proportion in the high-entropy intervals.
(b) Distributions on entropy, test loss, and test accuracy are analyzed. We randomly sample a dataset from our UIL scenario and categorize the samples into entropy intervals. The results, including the ratio of each interval, along with the average loss and average accuracy during the inference phase, indicate that lower entropy lead to higher accuracy.
Figure 2: Analysis on entropy distribution in the proposed UIL scenario, as well as the VIL scenario.
3.2 Challenge Analysis
To analyze the challenges of our UIL, we perform a detailed investigation. At the inter-task distribution level, we individually calculate PDE of all test samples in the UIL and VIL scenario after incremental training on the same $T=30$ tasks using the dataset [27] with various classes and domains. Then we divide all test samples into multiple intervals based on $H(y_{i})$ in ascending order, such as =$\mathcal{I}=\{\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3},\cdots,\mathcal{I}_{27}\}$, and count the ratio of samples in each interval to all samples. As illustrated in Fig.2(a), the UIL scenario results in more confusion than the VIL scenario, as proved by two key observations: 1) The ratio of samples in the low-entropy intervals is significantly smaller. 2) The ratio of samples in the high-entropy intervals is relatively larger. Besides, Fig. 2(b) indicates that lower entropy samples lead to more accurate predictions, suggesting that reducing entropy distribution is an effective way to improve model accuracy.
At the intra-task distribution level, during training phase, we calculate the average gradient magnitudes of the corresponding class’s weight vectors across three random experiments. In addition, the average test accuracy of corresponding classes is computed during inference phase. The joint relationship among the classes, gradient magnitude and test accuracy is illustrated in Fig.3. We observe a positive trend that the larger gradient magnitude is significantly associated with higher accuracy. This causes the model to learn knowledge from head classes in a biased manner if no any intervention is given, referred to asymmetrical optimization. Consequently, the model is confused about tail classes and biased to predict head classes, failing to achieve comprehensive performance across classes.
Figure 3: Illustration on the joint relationship among classes, corresponding gradient magnitude and test accuracy in a 3D bar chart. The imbalanced class distribution are resorted in ascending order based on the number of samples. This figure indicates that there is positive trend where higher gradient magnitudes correlate with improved accuracy, offering valuable insights into the performance across different classes.
Figure 4: Overview of our proposed UIL scenario (left) and our framework MiCo (right). During training on $\mathcal{D}_{train}=\mathcal{D}_{t}$, MiCo employs a multi-objective learning scheme and introduces direction- and magnitude-decoupled recalibration modules to mitigate confusion from inter- and intra-task distribution randomness. $\mathcal{L}_{ce}$ and $\mathcal{L}_{em}$ jointly enforce the model to make accurate and deterministic predictions with the role of reducing conflict between $\vec{g}_{ce}$ and $\vec{g}_{em}$) via a direction recalibration module. Moreover, a magnitude recalibration module is used to rescale gradient magnitude (e.g., $||g_{ce}||$, $||g_{em}||$), aiming to alleviate asymmetrical optimization towards imbalanced class distribution. Finally, gradient $\hat{g}_{f}$ with recalibrated direction $\vec{g}_{f}$ and magnitude $||\hat{g}_{f}||$ is used to update parameters in backward propagation. During inference, all previous tasks $\mathcal{D}_{test}=\{\mathcal{D}_{1},\cdots,\mathcal{D}_{t}\}$ are evaluated.
4 Method
Inspired by these observations in Sec. 3, we propose a simple yet effective framework for the UIL scenario, named MiCo. As illustrated in Fig. 4, the architecture of MiCo consists of a frozen pre-trained model and a trainable classifier during training on $\mathcal{D}_{train}=\mathcal{D}_{t}$, and all parameters are frozen during evaluation on previous tasks $\mathcal{D}_{test}=\{\mathcal{D}_{1},\cdots,\mathcal{D}_{t}\}$. In Sec. 4.1, to mitigate confusion from inter-task distribution randomness, we firstly employ a multi-objective learning scheme, referred to joint training using $\mathcal{L}_{ce}$ and $\mathcal{L}_{em}$, to enforce the model to produce accurate and deterministic predictions for random IL types across all task distributions. Furthermore, we introduce a direction recalibration module to enhance its effectiveness by reducing conflicting gradients, referred to $\vec{g}_{ce}=\frac{g_{ce}}{||g_{ce}||}$ and $\vec{g}_{em}=\frac{g_{em}}{||g_{em}||}$. In Sec. 4.2, we introduce how to mitigate confusion from intra-task distribution randomness. Towards training on imbalanced class distribution, we introduce a magnitude recalibration module to alleviate the challenge of asymmetrical optimization via rescaling gradient magnitude, referred to $||g_{ce}||$ and $||g_{em}||$ The gradient $\hat{g}_{f}$ after recalibration is finally used to update parameters in backward propagation.
4.1 Mitigating Confusion from Inter-task Distribution Randomness
In the CIL or DIL scenario, the incremental tasks have class- or domain- priority. Specifically, only the same number of new classes or domains, are sequentially introduced, allowing the model to confidently learn class- or domain-specific knowledge. However, there is no prior knowledge of the incremental type in the proposed UIL scenario. As a result, it confuses the model that the new knowledge introduced by new incremental types conflicts with that it has already learned, leading to confused prediction towards previous tasks. To address this, we employ a multi-objective learning scheme, which guide the model to make accurate and consistent predictions in confused IL scenarios.
As shown in Fig. 4, we joint train the model using the cross-entropy loss $\mathcal{L}_{ce}$ and the entropy minimization (EM) loss $\mathcal{L}_{em}$, which is commonly used in test-time adaptation [35, 25]. Specifically, we treat $p(\hat{y}_{i,c})$ as the probability that the input sample $x_{i}$ is classified as $c$-th label, where $\hat{y}_{i,c}$ is logits at the $c$-th position of the model’s output. The prediction distribution entropy (PDE) of $x_{i}$ is $H(y_{i})$, as shown in Eq. 2. Across all task distributions with random incremental types, given to the sample $x_{i}$ with the random label $y_{i}$, we guide the model to make accurate and deterministic predictions by teaching to correctly distinguish groud-truth label and explicitly reducing its PDE. Overall, MiCo is trained using the following multi-objective loss $\mathcal{L}$, as shown in Eq.5:
$\mathcal{L}=\mathcal{L}_{ce}+\gamma\cdot\mathcal{L}_{em},$ (5)
where $\gamma$ is balanced coefficient, discussed in ablation study as shown in Table. 3.
To enhance the effectiveness of the multi-objective learning scheme, the direction recalibration module is introduced to reduce the conflicting gradients between $\vec{g}_{ce}$ and $\vec{g}_{em}$. This is based on the observation that the gradient cosine similarity (GCS) is negative. Specifically, denote $\mathcal{G}_{d}=\{\vec{g}^{1},\vec{g}^{2},\cdots,\vec{g}^{c}\}$ as the set of gradient direction, where $c$ individually the class index of $c$-th weight vector in the classifier. As shown in Eq. 6:
$\vec{g}^{c}$ $=\vec{g}_{ce}^{c}+\gamma\cdot\vec{g}_{em}^{c},$ (6)
$\vec{g}_{f}^{c}$ $=\vec{g}^{{c}}+\beta\cdot\vec{g}_{w}^{c},$
where $\gamma$ is the same coefficient as Eq. 5 and $\vec{g}^{c}$ is original gradient direction before the direction recalibration module. $\beta=f(||\vec{g}^{c}||)$ is constant to $c$-th index and $\vec{g}_{w}^{c}$ is the recalibration offset, which can be solved in dual optimization, referred to [18]. Finally, $\vec{g}_{f}^{c}$ is the gradient direction of classifier’s $c$-th weight vector after the direction recalibration module. As a result, with the multi-objective learning scheme and the avoidance of conflicting gradient in the direction module, our model achieves not only accurate and deterministic predictions but also learns in a conflict-free direction, regardless of the variation in incremental types.
4.2 Mitigating Confusion from Intra-task Distribution Randomness
In the dynamic wild, the scale of these increments (i.e. the number of classes) is random, so the class distribution is eventually imbalanced due to distinct emergent frequencies. The standard optimization algorithm [16, 14] is to calculate the average loss across samples in a mini-batch and update parameters accordingly, regardless of whether class distribution is balanced or not. We identify a major challenge towards training on imbalanced class distribution, named asymmetrical optimization, where the classifier’s weight vectors corresponding to head classes experience substantial updates due to larger gradient magnitudes than tail classes.
To address this, we introduce a magnitude recalibration module to alleviate the asymmetrical optimization by rescaling the classifier’s gradient magnitude. Specifically, denote $\mathcal{G}_{m}=\{||g^{1}||,||g^{2}||,\cdots,||g^{c}||\}$ as the set of gradient magnitude, where $c$ is the class index of $c$-th weight vector. $g^{c}$ is defined in Eq. 7:
$||g^{c}||=||g^{c}_{ce}||+\gamma\cdot||g_{em}^{c}||,$ (7)
where $||g^{c}_{ce}||$ , $||g_{em}^{c}||$ and $||g^{c}||$ are gradient magnitude from $\mathcal{L}_{ce}$, $\mathcal{L}_{em}$, and total loss $\mathcal{L}$, respectively. $\gamma$ is the same balanced coefficient as Eq.5. We rescale each gradient magnitude in $\mathcal{G}_{m}$ using class-specific factor ${w}$ derived from Min-Max Normalization on $\mathcal{G}_{m}$. As shown in Eq. 8:
$||\hat{g}^{i}||$ $=w_{i}*||g^{i}||,$ (8)
where $\text{where }w_{i}$ $=\frac{||g^{i}||-\min(\mathcal{G}_{m})}{\max(\mathcal{G}_{m})-\min(\mathcal{G}_{m})}.$
Finally, the recalibrated gradient $\hat{g}^{c}_{f}$ with direction $\vec{g}^{c}_{f}$ and magnitude $||\hat{g}^{c}_{f}||$, as shown in Eq.6 and Eq.8, is used to update the classifier’s $c$-th weight in backward propagation. With the magnitude calibration module, our model effectively alleviates asymmetrical optimization by simple rescaling strategy when faced with imbalanced class distribution, allowing it to fairly acquire knowledge without learning bias, thereby achieving comprehensive performance across all classes.
5 Experiments
In this section, we conduct extensive experiments for comparison with state-of-the-art methods. First, we describe the experiment setup including datasets, evaluation metrics and comparison methods in Sec. 5.1. Moreover, we present the experimental results to demonstrate the superiority of our approach in Sec. 5.2. Finally, the effectiveness of each component is verified through ablative studies in Sec. 5.3.
5.1 Experiment Setup
Datasets and Metrics. Following the previous methods [26, 39, 7, 23], our experiments are conducted on datasets with various classes and domains: iDigits [34], CORe50 [19], DomainNet [27], which can easily be investigated in the realistic UIL and VIL scenario. The detailed task configuration is described in Table. 1. As previous evaluation protocol in [39, 33, 38, 26], we report the performance after training the last task $T$ using commonly evaluation metrics: average accuracy (marked as Avg. Acc) and forgetting rate (marked as Forgetting). Besides, due to the randomness causing an imbalanced data distribution across task distributions, we also report the weighted accuracy (marked as W. Acc) on all tasks for a comprehensive evaluation, where each task’s accuracy is averaged by the weight of each task.
Comparison Methods and Implementation Details. We compare MiCo with the state-of-the-art IL methods of three categories: prompt-based methods [39, 38, 33, 37], regularization-based methods [26, 8], mixture-based methods [7]. In addition, we also conduct the fine-tuning and joint-training experiments to provide the lower bound and upper bound of performance in a linear probe way [28, 9, 10]. All IL methods uniformly use ViT-B/16 [4] as the pre-trained backbone. Each method is trained over three random experiments with different seeds, and we report final performance as the mean and standard deviation. For training configuration (i.e., learning rate, epochs), we keep the same setting as ICON [26].
Table 1: Task configuration on three benchmarks in the UIL and VIL scenario. $T,\mathcal{C}_{t}$ and $\mathcal{M}_{t}$ denote the total number of sequential tasks, the number of classes and the number of domains within each incremental task, respectively. In the UIL scenario, $unk$ indicates that the number of classes and domains are unknown due to random scale of these increments. The value with asterisk (*) indicates that we try to ensure that each task covers multiple domains, preventing the UIL scenario from degrading into the VIL scenario for iDigits with fewer classes and domains, which leads to fewer tasks overall.
Dataset VIL UIL
$T$ $\mathcal{C}_{t}$ $\mathcal{M}_{t}$ $T$ $\mathcal{C}_{t}$ $\mathcal{M}_{t}$
iDigits [34] 20 2 1 5* $unk$ $unk$
CORe50 [19] 40 10 1 40 $unk$ $unk$
DomainNet [27] 30 10 1 30 $unk$ $unk$
Table 2: Comparison with state-of-the-art IL methods on the UIL and VIL scenario. The comparison is shown in terms of Avg. Acc($\uparrow$), W. Acc($\uparrow$), and Forgetting($\downarrow$) as the mean and standard deviation across three random experiments. The Avg. Acc($\uparrow$) on both scenarios is averaged as the Mean Acc($\uparrow$) at the last column. Since joint training, as the upper-bound performance, is inherently scenario-agnostic for incremental learning, so the only Avg. Acc($\uparrow$) is reported. For a comprehensive comparison, we compare MiCo with recently existing methods of prompt-based [39, 33, 38], mixture-based [7] and regularization-based [26, 8]. We maintain a fixed amount of memory with 10 samples per classes for replaying experiences in [8]. The best and second results are marked in the bold and underline, respectively.
Method VIL UIL Mean Acc($\uparrow$)
Avg. Acc($\uparrow$) W. Acc($\uparrow$) Forgetting($\downarrow$) Avg. Acc($\uparrow$) W. Acc($\uparrow$) Forgetting($\downarrow$)
iDigits
Upper-bound - 95.6$\pm$0.4
Lower-bound 36.3$\pm$0.1 32.6$\pm$1.0 21.5$\pm$5.5 46.5$\pm$0.5 50$\pm$2.4 18.4$\pm$13.0 41.4$\pm$0.3
L2P [39] 57.2$\pm$3.0 52.1$\pm$3.7 32.1$\pm$4.1 74.3$\pm$1.3 74.1$\pm$3.0 18.0$\pm$9.1 65.8$\pm$2.2
DualPrompt [38] 61.8$\pm$1.5 59.0$\pm$2.9 24.9$\pm$3.6 74.3$\pm$0.1 75.4$\pm$0.4 15.4$\pm$4.5 68.1$\pm$0.8
CODA-Prompt [33] 64.6$\pm$1.8 61.0$\pm$3.5 23.6$\pm$1.7 79.0$\pm$11.2 78.0$\pm$12.7 18.3$\pm$17.0 71.8$\pm$6.5
DGR [8] 65.3$\pm$3.2 63.9$\pm$2.6 38.0$\pm$3.7 79.3$\pm$4.9 77.6$\pm$8.1 25.4$\pm$5.3 72.3$\pm$4.1
LAE [7] 50.1$\pm$4.3 42.5$\pm$5.4 36.1$\pm$1.0 74.5$\pm$2.2 74.4$\pm$1.6 21.7$\pm$2.0 62.3$\pm$3.3
ICON [26] 68.0$\pm$2.0 65.1$\pm$2.5 21.3$\pm$2.7 81.8$\pm$8.0 80.2$\pm$10.3 15.7$\pm$13.1 74.9$\pm$5.0
Ours: MiCo 69.5$\pm$3.2 67.3$\pm$3.5 23.4$\pm$5.4 84.6$\pm$4.1 83.9$\pm$6.2 11.5$\pm$4.3 77.1$\pm$3.6
CORe50
Upper-bound - 95.3$\pm$0.3
Lower-bound 69.6$\pm$0.3 69.6$\pm$0.4 5.9$\pm$0.8 45.1$\pm$4.0 46.4$\pm$4.4 16.8 $\pm$0.2 57.4$\pm$2.2
L2P [39] 75.5$\pm$3.5 75.6$\pm$3.5 8.0$\pm$1.9 50.0$\pm$1.6 51.6$\pm$0.9 18.6$\pm$2.9 62.8$\pm$2.6
DualPrompt [38] 73.1$\pm$7.6 73.1$\pm$7.6 9.8$\pm$4.3 51.4$\pm$1.6 52.6$\pm$1.0 18.1$\pm$3.0 62.3$\pm$4.6
CODA-Prompt [33] 73.2$\pm$0.1 73.3$\pm$0.2 8.0$\pm$2.0 53.9$\pm$2.3 54.2$\pm$1.5 15.7$\pm$5.4 63.6$\pm$1.2
DGR [8] 77.5$\pm$3.3 77.0$\pm$2.9 12.4$\pm$3.8 55.8$\pm$3.6 56.7$\pm$2.4 20.1$\pm$3.5 66.7$\pm$3.5
LAE [7] 72.0$\pm$3.2 73.2$\pm$2.6 8.3$\pm$2.7 54.1$\pm$5.0 55.5$\pm$5.4 16.6$\pm$1.0 63.1$\pm$4.1
ICON [26] 80.6$\pm$1.1 80.7$\pm$1.0 6.2$\pm$0.1 57.5$\pm$1.2 59.1$\pm$1.1 15.8$\pm$4.4 69.1$\pm$1.2
Ours: MiCo 83.5$\pm$0.8 82.0$\pm$0.8 5.3$\pm$0.5 67.4$\pm$2.3 66.4$\pm$2.0 16.4$\pm$3.2 75.5$\pm$1.5
DomainNet
Upper-bound - 59.5$\pm$0.4
Lower-bound 42.0$\pm$2.8 43.4$\pm$1.5 15.2$\pm$1.9 42.1$\pm$3.6 45.9$\pm$0.2 8.3$\pm$2.8 42.1$\pm$3.2
L2P [39] 47.0$\pm$1.7 47.3$\pm$1.2 24.2$\pm$1.7 43.2$\pm$0.9 41.6$\pm$0.7 22.4$\pm$4.9 45.1$\pm$1.3
DualPrompt [38] 45.8$\pm$2.9 47.1$\pm$1.2 18.6$\pm$2.7 46.3$\pm$3.4 42.0$\pm$1.9 23.5$\pm$5.0 46.1$\pm$3.2
CODA-Prompt [33] 47.6$\pm$1.9 48.4$\pm$0.8 19.9$\pm$1.7 46.0$\pm$2.4 43.2$\pm$1.9 24.2$\pm$3.5 46.8$\pm$2.2
DGR [8] 48.3$\pm$1.5 48.8$\pm$1.4 19.6$\pm$2.2 49.2$\pm$3.0 50.1$\pm$1.8 23.4$\pm$3.8 48.8$\pm$2.3
LAE [7] 44.6$\pm$0.5 43.8$\pm$1.2 25.0$\pm$2.3 47.5$\pm$3.1 42.1$\pm$1.9 19.3$\pm$4.9 46.1$\pm$1.8
ICON [26] 49.8$\pm$0.3 52.6$\pm$1.7 14.4$\pm$1.3 52.3$\pm$2.6 53.0$\pm$1.5 15.2$\pm$2.6 51.1$\pm$1.5
Ours: MiCo 52.3$\pm$1.2 56.7$\pm$1.3 16.7$\pm$1.5 55.2$\pm$2.3 56.1$\pm$0.1 15.0$\pm$2.8 53.8$\pm$1.8
Table 3: Ablation experiments of each component in UIL. mobj is the multi-objective learning scheme. dir is the direction recalibration module. mag is the magnitude recalibration module. $\checkmark$ indicates the component is enables. We conduct ablative experiments by progressively integrating each component into the baseline from baseline (1) to MiCo (6). dir only exists in conjunction with mobj, aiming to reduce conflicting gradients. The Avg. Acc(↑) across three benchmarks is averaged as the Mean Acc(↑) at the last column.
ID Method iDigits CORe50 DomainNet Mean Acc($\uparrow$)
mobj dir mag Avg. Acc($\uparrow$) W. Acc($\uparrow$) Forgetting($\downarrow$) Avg. Acc($\uparrow$) W. Acc($\uparrow$) Forgetting($\downarrow$) Avg. Acc($\uparrow$) W. Acc($\uparrow$) Forgetting($\downarrow$)
(1) 78.3$\pm$3.2 79.5$\pm$6.1 14.3$\pm$7.9 64.9$\pm$5.6 64.7$\pm$6.9 20.5$\pm$ 4.3 46.9$\pm$3.8 55.2$\pm$1.4 11.6$\pm$2.6 63.3$\pm$4.2
(2) 80.4$\pm$4.6 81.0$\pm$5.9 13.5$\pm$9.5 65.0$\pm$3.2 65.5$\pm$4.4 19.7$\pm$2.3 49.8$\pm$3.0 55.7$\pm$1.4 16.6$\pm$1.4 65.1$\pm$3.6
(3) 81.3$\pm$4.4 82.4$\pm$6.3 13.0$\pm$9.1 65.9$\pm$3.8 65.9$\pm$4.6 18.3$\pm$4.2 50.2$\pm$3.1 54.5$\pm$1.4 14.9$\pm$1.9 65.8$\pm$3.7
(4) 84.0$\pm$3.0 83.9$\pm$4.7 12.4$\pm$6.9 67.1$\pm$2.5 66.9$\pm$2.1 17.1$\pm$2.8 53.1$\pm$2.6 55.9$\pm$0.3 14.5$\pm$0.7 68.1$\pm$2.7
(5) 83.1$\pm$3.9 82.5$\pm$6.0 13.4$\pm$7.4 66.4$\pm$1.2 67.3$\pm$2.3 19.0$\pm$2.3 51.2$\pm$3.1 55.2$\pm$0.3 15.8$\pm$5.1 67.0$\pm$2.7
(6) 84.6$\pm$4.1 83.9$\pm$6.2 11.5$\pm$4.3 67.4$\pm$2.3 66.4$\pm$2.0 16.4$\pm$3.2 55.2$\pm$2.3 56.1$\pm$0.1 15.0$\pm$2.8 69.1$\pm$2.9
Table 4: Ablation of $\gamma$ in UIL. The Avg. Acc(↑) across three benchmarks is averaged as the Mean Acc(↑) at the last column.
$\gamma$ iDigits CORe50 DomainNet Mean Acc($\uparrow$)
Avg. Acc($\uparrow$) W. Acc($\uparrow$) Forgetting($\downarrow$) Avg. Acc($\uparrow$) W. Acc($\uparrow$) Forgetting($\downarrow$) Avg. Acc($\uparrow$) Avg. Acc($\uparrow$) Forgetting($\downarrow$)
1.0 85.1$\pm$4.3 84.0$\pm$ 5.3 12.7$\pm$7.9 66.3$\pm$3.4 66.6$\pm$2.8 17.2$\pm$2.1 47.3$\pm$3.9 54.4$\pm$0.5 11.1$\pm$1.7 66.2$\pm$3.9
0.1 80.5$\pm$7.9 80.8$\pm$ 8.4 15.8$\pm$10.2 66.5$\pm$4.5 66.4$\pm$5.6 17.8$\pm$5.2 52.0$\pm$1.2 55.5$\pm$0.7 15.3$\pm$4.9 66.3$\pm$4.5
0.01 84.6$\pm$4.1 83.9$\pm$6.2 11.5$\pm$4.3 67.4$\pm$2.3 66.4$\pm$2.0 16.4$\pm$3.2 55.2$\pm$2.3 56.1$\pm$0.1 15.0$\pm$2.8 69.1$\pm$2.9
0.001 83.2$\pm$2.1 82.1$\pm$5.8 13.6$\pm$7.6 66.5$\pm$2.1 66.2$\pm$0.3 22.0$\pm$0.6 52.0$\pm$0.1 54.9$\pm$0.1 19.1$\pm$1.3 67.2$\pm$1.4
5.2 Comparison with State-of-The-Art IL Methods
We compare MiCo with the existing method in realistic IL scenarios, including the proposed UIL scenario and the VIL scenario. As shown in Table. 2, MiCo significantly outperforms existing sota methods in general UIL and VIL scenario in terms of all evaluation metrics. Compared with second results of 74.9, 69.1, 51.1 in terms of Mean Acc ($\uparrow$) on three benchmarks, our MiCo achieves significant gains with 77.1 (+2.2), 75.5 (+6.4), 53.8 (+2.7) and simultaneously has comparable variance with 3.6 (-1.4), 1.5 (+0.3), 1.8 (+0.3) in the corresponding benchmarks. Moreover, the ability of MiCo to overcome catastrophic forgetting also maintains a considerable mean Forgetting ($\downarrow$) with the second place [26]. These observations demonstrate that MiCo has a more robust ability to mitigate confusion from inter- and intra-task distribution randomness without degrading the ability to overcome forgetting. Existing prompt-based methods L2P[39], DualPrompt [38], CODA-Propmt [33] aim to select generalizable prompts by utilizing a cosine similarity-based matching mechanism to mitigate catastrophic forgetting without storing previous rehearsal. It is ineffective when domain variation continually occurs within classes in the UIL scenario, because similarity-based matching is extremely sensitive to appearance variation. Among regularization-based methods, ICON [26] introduces specific regularization terms by enforcing orthogonality among adapter shifts to alleviate forgetting but shows suboptimal performance due to random incremental scales. DGR [8] relies on the prior distribution of the stored rehearsals to distill the previous knowledge but gets confused about varying incremental types. LAE [7], a representative mixture-based method, continually merges online and offline models into one unified model to overcome forgetting. However, its performance is not unstable in the more confused IL scenario. Instead, the results in Table. 2 indicates that our MiCo achieves reliable performance in both scenarios. Actually, our UIL inevitably involves knowledge overlap across all tasks (i.e. the former and the latter tasks share the same classes or domains). Therefore, vanilla fine-tuning even without any specific architectures can performs well.
5.3 Ablation Study
Effectiveness of Components. As shown in Table. 3, we conduct ablative experiments by progressively integrating each component into our baseline to verify the effectiveness of each component. Compared with Mean Acc ($\uparrow$) of 63.3$\pm$4.2 of baseline (1), the effectiveness of the multi-objective learning scheme in (2) is validated, achieving significant gains with 65.1$\pm$3.6 and surpassing the baseline by 1.8 points. The results demonstrate that the multi-objective learning scheme contributes to making accurate and deterministic predictions regardless of random incremental types across all task distributions. Additionally, the effectiveness of the magnitude recalibration module in (3) is verified, achieving significant gains with 65.8$\pm$3.7 and surpass the baseline by 2.5 points. The results demonstrate that the magnitude recalibration module could effectively alleviate asymmetrical optimization towards imbalanced class distribution. By simultaneously adopting the multi-objective learning scheme and the magnitude recalibration module in (4), the effectiveness of each component can be leveraged through synergy, achieving absolute gains up to 3.0 and 2.3 points, respectively. To ablate the direction recalibration module, we combine the multi-objective learning with the direction recalibration module in (5) and (6), respectively. The effectiveness of the multi-objective learning scheme is further enhanced. Although our direction calibration module indirectly modifies the gradient magnitude, there is synergistic when simultaneously enabling direction- and magnitude-decoupled recalibration modules, indicated by the comparison of (4), (5) and (6). Overall, each component of our MiCo (6) is cooperative and effective in improving accuracy and overcoming forgetting for the UIL scenario, outperforming the baseline (1) with accuracy gains up to 5.8 points and forgetting gains down to 1.2 points.
Ablation of $\gamma$. In our multi-objective learning scheme, we perform three random experiments with different seeds to find the best $\gamma$, i.e. $\gamma\in\{1.0,0.1,0.01,0.001\}$. As shown in Table.4, when $\gamma$ set to 0.01, it exhibits the best performance at the peak on both CORe50 [19] and DomainNet [27], particularly in terms of Avg. Acc($\uparrow$). Although $\gamma=0.01$ fails to achieve the best accuracy on iDigits [34], it has comparable performance in terms of accuracy and the best performance in terms of forgetting. The results indicate that excessively larger values can lead to degrading accuracy, while excessively small values are unable to make deterministic predictions. Therefore, $\gamma$ is set to 0.01 by default in our experiments.
6 Conclusion
In this paper, we investigate a more general and realistic Universal Incremental Learning (UIL), where the model has no prior knowledge of the types and scale of increments. Furthermore, we conduct the detailed analysis for the UIL scenario and identify confusion arising from inter- and intra-task distribution randomness. To mitigate confusion, we propose a simple yet effective framework named MiCo, consisting of a multi-objective learning scheme and direction- and magnitude-decoupled recalibration modules. Extensive experiments demonstrate that our MiCo shows sota performance on the UIL scenario, and existing VIL scenario. We hope that UIL could provide new insights for IL community. In the future, our focus will be on establishing a comprehensive benchmark for UIL to accurately measure performance.
Acknowledgement
This work was supported by the National Natural Science Foundation of China (Nos.62476054, and 62172228).

### v2

@@
 Figure 1: Comparison of existing IL scenarios and our UIL scenario on a 6-class dataset with 4 domains.
-The horizontal axis represents class variation, while the vertical axis represents domain variation.
-Each incremental task is framed by a black box.
-UIL is a more general IL scenario than existing IL scenarios by simultaneously considering randomness in both the type and scale of increments, as the model neither knows which new classes or domains will increase across all task distributions, nor the number of new classes or domains within each task distribution.
+The horizontal axis denotes class index and the vertical axis denotes domain index.
+Each incremental task is outlined in black.
+CIL, DIL, and VIL assume that each incoming task introduces only class increments, domain increments, or both class and domain increments, respectively.
+In contrast, UIL is a more general IL scenario than existing ones by simultaneously introducing randomness in both the type and scale of increments, as the model neither knows which new classes or domains will increase across all task distributions, nor the number of new classes or domains within each task distribution.
 1 Introduction
@@
 Specifically, denote $\mathcal{G}_{m}=\{||g^{1}||,||g^{2}||,\cdots,||g^{c}||\}$ as the set of gradient magnitude, where $c$ is the class index of $c$-th weight vector.
-$g^{c}$ is defined in Eq.
+$||g^{c}||$ is defined in Eq.
 7:
-$||g^{c}||=||g^{c}_{ce}||+\gamma\cdot||g_{em}^{c}||,$ (7)
-where $||g^{c}_{ce}||$ , $||g_{em}^{c}||$ and $||g^{c}||$ are gradient magnitude from $\mathcal{L}_{ce}$, $\mathcal{L}_{em}$, and total loss $\mathcal{L}$, respectively.
+$||g^{c}||=||g^{c}_{ce}+\gamma\cdot g_{em}^{c}||,$ (7)
+where $g^{c}_{ce}$ and $g_{em}^{c}$ are gradient vector from $\mathcal{L}_{ce}$ and $\mathcal{L}_{em}$, respectively.
+$||g^{c}||$ is the accumulated gradient magnitude after vector summation.
 $\gamma$ is the same balanced coefficient as Eq.5.
@@
 8:
-$||\hat{g}^{i}||$ $=w_{i}*||g^{i}||,$ (8)
-where $\text{where }w_{i}$ $=\frac{||g^{i}||-\min(\mathcal{G}_{m})}{\max(\mathcal{G}_{m})-\min(\mathcal{G}_{m})}.$
+$||\hat{g}^{c}||$ $=w_{c}*||g^{c}||,$ (8)
+where $\text{where }w_{c}$ $=1-\frac{||g^{c}||-\min(\mathcal{G}_{m})}{\max(\mathcal{G}_{m})-\min(\mathcal{G}_{m})}.$
 Finally, the recalibrated gradient $\hat{g}^{c}_{f}$ with direction $\vec{g}^{c}_{f}$ and magnitude $||\hat{g}^{c}_{f}||$, as shown in Eq.6 and Eq.8, is used to update the classifier’s $c$-th weight in backward propagation.
@@
 Each method is trained over three random experiments with different seeds, and we report final performance as the mean and standard deviation.
-For training configuration (i.e., learning rate, epochs), we keep the same setting as ICON [26].
+For training configuration (i.e., learning rate, epochs and learnable adapter for fine-tuning backbone), we keep the same setting as ICON [26].
 Table 1: Task configuration on three benchmarks in the UIL and VIL scenario.
-$T,\mathcal{C}_{t}$ and $\mathcal{M}_{t}$ denote the total number of sequential tasks, the number of classes and the number of domains within each incremental task, respectively.
+$T$,$\left\|\mathcal{C}_{t}\right\|$ and $\left\|\mathcal{M}_{t}\right\|$ denote the total number of sequential tasks, the number of classes and the number of domains within each incremental task, respectively.
 In the UIL scenario, $unk$ indicates that the number of classes and domains are unknown due to random scale of these increments.
@@
 Dataset VIL UIL
-$T$ $\mathcal{C}_{t}$ $\mathcal{M}_{t}$ $T$ $\mathcal{C}_{t}$ $\mathcal{M}_{t}$
+$T$ $\left\|\mathcal{C}_{t}\right\|$ $\left\|\mathcal{M}_{t}\right\|$ $T$ $\left\|\mathcal{C}_{t}\right\|$ $\left\|\mathcal{M}_{t}\right\|$
 iDigits [34] 20 2 1 5* $unk$ $unk$